OpenStudy (anonymous):

Two large Fibonacci numbers are F51 = 20365011074 and F52 = 32951280099. If these two numbers are added, what do you get a f52 b f53 c f51 d f49 e f50

Vocaloid (vocaloid):

remember the rule for the Fibonacci sequence: each term is the sum of the previous two terms F(N+2) = F(N) + F(N+1) let N = 51

OpenStudy (anonymous):

so f54

Vocaloid (vocaloid):

no, please check your math again.

Vocaloid (vocaloid):

F(N+2) = ?